# Fast Training of Sinusoidal Neural Fields via Scaling Initialization

### v1

@@ C.3 Non-kernel approach: analysis via Hessian eigenvalues
Understanding the structure of the Hessian is a crucial aspect in the field of optimization. The Hessian matrix represents the second-order derivatives of the loss function with respect to each parameter, and its eigenvalues provide insights into the local curvature of the model’s parameter space. The condition number of the Hessian, defined as $\lambda_{\max}/\lambda_{\min}$, directly affects the convergence rate and guarantees of quadratic optimization models. In deep learning, the Hessian is widely studied to better understand the complex optimization trajectories of deep neural networks. For instance, Ghorbani et al. (2019) investigated outlier eigenvalues of the Hessian in neural networks (i.e., eigenvalues detached from the main bulk) and demonstrated through empirical and theoretical analysis that such outliers can slow down the convergence of the network.
@@
Building on this line of research, in this section, we present the condition number (CN) of the Hessian during training in three different settings. Fig. 10 illustrates that the CN [^2] For numerical stability, similar to Section 4.3, we compute the average of the top 3 eigenvalues for $\lambda_{\max}$ and the bottom 3 eigenvalues for $\lambda_{\min}$, where the scale of the eigenvalues are considered in terms of their absolute values. For Hessian computation, we use ‘PyHessian’ (Yao et al., 2020). of WS-SNF remains significantly lower than that of other comparison models throughout training. This suggests that WS initialization leads to a better-conditioned Hessian not only in the kernel-approximated regime but also in the nonlinear regime, contributing to more efficient training dynamics.
@@ Appendix D Frequency spectrum of 2D images
Fig. 4 presents a comparative analysis of the frequency spectrum between weight-scaled SNF and 1D audio signals. To further validate our findings, we extend our analysis to Kodak image dataset. Following the methodology of Shi et al. (2022), we partition the frequency map into several discrete bands and compute the mean values within each band.
@@
Appendix E Experimental details and qualitative results
E.1 Training settings and baselines
@@
In this section, we provide detailed information about our experiments. For data fitting tasks, we used the Adam optimizer (Kingma, 2015) with a learning rate of 0.0001, without any learning rate scheduler, except for the occupancy field experiments (in which case, we use PyTorch learning rate scheduler). Each experiment was conducted with 5 different seeds, and we report both the average and standard deviation of the evaluation metric. We note that all baselines in this work are MLP architectures, with slight modifications to the input features or activation functions. We used NVIDIA RTX 3090/4090/A5000/A6000 GPUs for all experiments. Additionally, we set the hyperparameters of each network using a grid search, as shown in LABEL:tab:details.
@@
GaussNet (Ramasinghe & Lucey, 2022): For this architecture, we use the Gaussian activation function, i.e., $\sigma(Z)=\exp\left(-(s\cdot Z)^{2}\right)$, where $Z$ denotes the pre-activation matrix and $s$ is a hyperparameter that defines the bandwidth of the Gaussian function.
@@
Neural field scaling law (Saratchandran et al., 2024): We modify the initialization distribution of the last layer of SIREN, i.e., $W^{(L)}\sim\mathcal{U}\left(-\sqrt{6}/\left(n^{3/4}_{L-1}\omega\right),\sqrt{6}/\left(n^{3/4}_{L-1}\omega\right)\right)$, where $n_{L-1}$ denotes the width of the $(L-1)$-th layer.
Xavier uniform initialization (Glorot & Bengio, 2010): In this case, we use the standard Xavier uniform initialization for the SNF architecture, i.e., $W^{(l)}\sim\mathcal{U}\left(-\sqrt{6}/\sqrt{n_{l-1}+n_{l}},\sqrt{6}/\sqrt{n_{l-1}+n_{l}}\right),\quad l\in[1,L]$.
@@
E.2 Neural dataset experiements
@@
SIREN 35 14m 23s 97.95$\pm$0.00 19m 14s 47.98$\pm$0.01
40 18m 20s 97.76$\pm$0.00 29m 45s 48.46$\pm$0.01
45 23m 23s 97.53$\pm$0.01 47m 54s 48.09$\pm$0.01
WS ($\times$ 1.5) 35 7m 22s 97.56$\pm$0.00 10m 52s 47.15$\pm$0.02
40 8m 22s 97.81$\pm$0.00 17m 21s 47.72$\pm$0.01
45 10m 41s 97.02$\pm$0.01 25m 52s 46.08$\pm$0.03
WS ($\times$ 2.0) 35 5m 25s 97.48$\pm$0.01 8m 26s 43.30$\pm$0.02
40 5m 32s 97.06$\pm$0.01 9m 20s 42.86$\pm$0.02
45 5m 41s 97.49$\pm$0.00 10m 51s 42.70$\pm$0.01
E.3 Training for additional iterations
Weight scaling initialization primarily focuses on the training efficiency of neural fields. In other words, the experiments in the main paper were conducted over a limited number of epochs. In this subsection, we provide qualitative results from models trained for additional iterations. As shown in Table 4, networks with WS initialization continue to outperform other networks, suggesting that its effectiveness is consistently preserved throughout the training process.
Image (300 iterations) Occ. Field (100 iterations) Spherical Data (15k iterations) Audio (5k iterations)
Activation KODAK DIV2K Lucy ERA5 Bach
Xavier Uniform (Glorot & Bengio, 2010) Sinusoidal 2.31 2.20 0.0706 7.10$\pm$1.02 16.16
ReLU + P.E. (Mildenhall et al., 2020) ReLU 21.77 19.60 0.9920 41.32$\pm$0.16 32.40
FFN (Tancik et al., 2020) ReLU 26.51 25.62 0.9874 45.27$\pm$0.12 17.54
SIREN init. (Sitzmann et al., 2020b) Sinusoidal 26.01 24.76 0.9926 44.42$\pm$0.32 46.78
GaussNet (Ramasinghe & Lucey, 2022) Gaussian 21.90 20.22 0.9935 44.15$\pm$0.12 34.65
WIRE (Saragadam et al., 2023) Wavelet 31.47 30.89 0.9939 45.13$\pm$0.42 20.17
NFSL (Saratchandran et al., 2024) Sinusoidal 26.91 25.97 0.9933 45.10$\pm$0.08 45.70
Weight scaling (ours) Sinusoidal 52.73 53.10 0.9950 49.63$\pm$0.14 54.04
Table 4: Quantitative results: training for additional iterations.
E.4 Qualitative results
@@
(a) Ground truth
(b) SNF + Xavier
(c) ReLU + P.E.
(d) FFN
(e) SIREN
(f) GaussNet
(g) WIRE
(h) NFSL
(i) WS (ours)
@@
(a) Ground truth
(b) ReLU + P.E.
(c) FFN
(d) SIREN
(e) GaussNet
(f) WIRE
(g) NFSL
(h) WS (ours)
@@
(a) Ground truth
(b) SNF + Xavier
(c) ReLU + P.E.
(d) FFN
(e) SIREN
(f) GaussNet
(g) WIRE
(h) NFSL
(i) WS (ours)
@@
(a) Ground truth
(b) SNF + Xavier
(c) ReLU + P.E.
(d) FFN
(e) SIREN
(f) GaussNet
(g) WIRE
(h) NFSL
(i) WS (ours)

### v3

@@ C.4 Non-kernel approach: analysis via Hessian eigenvalues
Understanding the structure of the Hessian is a crucial aspect in the field of optimization. The Hessian matrix represents the second-order derivatives of the loss function with respect to each parameter, and its eigenvalues provide insights into the local curvature of the model’s parameter space. The condition number of the Hessian, defined as $\lambda_{\max}/\lambda_{\min}$, directly affects the convergence rate in gradient-based convex optimization problems, which is provable (Nesterov and others, 2018).
On the other hand, in deep learning, the Hessian is also widely studied to gain a better understanding of the complex optimization trajectories and characteristics of deep neural networks. For instance, Ghorbani et al. (2019) examined the outlier eigenvalues of the Hessian in neural networks (i.e., eigenvalues detached from the main bulk) and demonstrated through comprehensive empirical analyses that these outliers can slow down the network’s convergence.
@@
Building on this line of research, in this subsection, we plot the condition number (CN) of the Hessian during training in three different settings. Fig. 10 illustrates that the CN [^3] For numerical stability, similar to Section 4.3, we compute the average of the top 3 eigenvalues for $\lambda_{\max}$ and the bottom 3 eigenvalues for $\lambda_{\min}$, where the scale of the eigenvalues are considered in terms of their absolute values. For Hessian computation, we use ‘PyHessian’ (Yao et al., 2020). of WS-SNF remains significantly lower than that of other comparison models throughout training. This suggests that WS initialization leads to a better-conditioned Hessian not only in the kernel-approximated regime but also in the nonlinear regime, contributing to more efficient training dynamics.
To provide deeper insights and establish connections with related work on spectral analyses of the Hessian in neural networks, we plot the eigenspectrum snapshot of the Hessian during training in Fig. 11. In both networks, we observe the occurrence of ‘outlier eigenvalues’ during training. However, the default setting exhibits a much more extreme value (even at initialization), which serves as a proxy for poorly conditioned optimization and slow convergence, aligning with the observations in Ghorbani et al. (2019) and Rathore et al. (2024).
@@
(a) Default (0 step)
(b) Default (600 step)
(c) Default (1339 step)
(d) Weight scaling (0 step)
(e) Weight scaling (35 step)
(f) Weight scaling (79 step)
Figure 11: Eigenspectrum of the Hessian during training: The first column represents the eigenspectrum at initialization, the second column corresponds to the eigenspectrum at mid-term training, and the third column shows the eigenspectrum at the end of training.
@@ Appendix D Frequency spectrum of 2D images
To further validate our findings, we extend our analysis to Kodak image dataset. Following the methodology of Shi et al. (2022), we partition the frequency map into several discrete bands and compute the mean values within each band.
@@
Appendix E More discussion
E.1 The effect of weight scaling in mini-batch training
Most of our primary experiments were conducted under a full-batch training to prioritize fast convergence. In this section, we provide supplementary experiments on mini-batch training for image regression to further investigate the effect of weight scaling.
Figure 13: PSNR curves for mini-batch training. PSNR curves for default and weight scaled SNFs with various batch sizes. Darker colors represent larger batch sizes.
We trained SNFs on a $\times512$ sized single image using five different batch sizes $\{2^{17},2^{16},2^{15},2^{14},2^{13}\}$ over 500 iterations. Fig. 13 demonstrates that weight scaling is also effective in mini-batch training, achieving significant acceleration compared to standard SNFs. These results emphasize the robustness and applicability of the proposed method.
E.2 On the effect of weight scaling on other activation functions
(a) SNF
(b) SNF
(c) ReLU + P.E.
(d) ReLU + P.E.
(e) GaussNet
(f) GaussNet
(g) WIRE
(h) WIRE
(i) Sinc NF
(j) Sinc NF
Figure 14: Weight scaling in various architectures. WS substantially affects training of the neural field with periodic activations. We present curves for PSNR (first column), as well as Test PSNR (second column).
We extend our investigation to examine the effect of weight scaling on different activation functions. Specifically, we evaluate ReLU, the most commonly used homogeneous activation function, along with other periodic activations beyond sinusoidal. Fig. 14 depicts the train and test PSNR curves over 500 iterations with varying values of $\alpha$. As expected, neural networks with ReLU activations are minimally affected by the scale of weight initialization. In contrast, architectures, such as GaussNet, WIRE, and Sinc NF are significantly influenced by weight scaling. This is because the weight distribution in non-homogeneous networks has a substantial impact on their performance, primarily through its effect on the functional frequency, as discussed in our manuscript.
We observe that weight-scaled GaussNet and Sinc NF also enjoy better performance in terms of the training PSNR, but their test PSNR degrades more severely than in SNFs. Moreover, WIRE shows a degraded performance with weight scaling. While our analysis primarily focused on sinusoidal neural fields, we emphasize that initialization for other periodic activations also impacts acceleration and generalization of neural field training. We leave this as future work, aiming to provide valuable insights into the role of weight initialization in enabling faster convergence.
E.3 The optimal scaling factor according to the learning rate
Figure 15: Optimal scaling factor vs. learning rate.
We discussed the relationship between the optimal scaling factor and structure properties of neural fields in Section 5.2. However, the learning rate significantly influences training dynamics, potentially leading to instability within the hyperparameter space. Fig. 15 illustrates the optimal weight scaling factor $\alpha$ for different learning rates $\eta\in\{10^{-3},10^{-4},10^{-5}\}$. Our observations indicate that weight scaling consistently accelerates training across all tested learning rates. Furthermore, a clear trend emerges: smaller values of $\alpha$ are optimal for larger $\eta$. This trend highlights the predictability of the optimal $\alpha$, consistent with other structural properties discussed in our main section.
Appendix F Experimental details and qualitative results
F.1 Experimental details about Fig. 2(b)
Fig. 2(b) presents the tradeoff between acceleration and generalization when tuning $\omega$ or $\alpha$. Therefore, we choose the ranges of $\omega$ and $\alpha$ that can help us capture the full tradeoff curve. Specifically, for frequency tuning, we explore a broad range, $\omega\in[30,6510]$, with intervals of 270, while also including smaller values ($\omega=10$ and $20$) in the search space. On the other hand, weight scaling requires a narrower search range, with $\alpha\in[0.4,40]$ with the grid $0.2$.
F.2 Experimental details about Fig. 3
For the $k$-th layer of an $l$-layer SNF, weight scaling increases the gradient magnitude by $\alpha^{l-k}$ compared to standard setting. By scaling down the learning rate of each layer by $\eta\cdot\alpha^{l-k}$, the gradient update magnitudes become similar to those of the default while preserving the initial functional of weight scaling (Fig. 3(b)). In contrast, for Fig. 3(c), the learning rate of $k$th layer is amplified by $\eta\cdot\alpha^{l-k}$ to only utilize the effect of larger gradients. This is inspired by our analytic derivations at the end of Section 4.1 (and Section B.3).
F.3 Experimental details about Eq. 10
We clarify that the optimization Eq. 10 is solved by a simple grid search. That is, we have trained SNFs with many different values of $\alpha$ and selected the value which maximizes the training speed and meets the constraint. We have tried all $\alpha$ within the range $[1.0,4.0]$ with the grid size $0.2$.
F.4 Training settings and baselines
@@
In this section, we provide detailed information about our experiments. For data fitting tasks, we used the Adam optimizer (Kingma, 2015) with a learning rate of 1e-04 (except spherical data, in this case we use a learning rate of 1e-05), without any learning rate scheduler, except for the occupancy field experiments (in which case, we use PyTorch learning rate scheduler). Each experiment was conducted with 5 different seeds, and we report both the average and standard deviation of the evaluation metric. We note that all baselines in this work are MLP architectures, with slight modifications to the input features or activation functions. We used NVIDIA RTX 3090/4090/A5000/A6000 GPUs for all experiments. Additionally, we set the hyperparameters of each network using a grid search, as shown in LABEL:tab:details.
@@
MFN (Fathony et al., 2021): For this architecture, we use MLP with multiple Gabor filters, which achieves superior results than Fourier filter networks (Fathony et al., 2021).
GaussNet (Ramasinghe and Lucey, 2022): For this architecture, we use the Gaussian activation function, i.e., $\sigma(Z)=\exp\left(-(s\cdot Z)^{2}\right)$, where $Z$ denotes the pre-activation matrix and $s$ is a hyperparameter that defines the bandwidth of the Gaussian function.
@@
Neural field scaling law (NFSL) (Saratchandran et al., 2024b): We modify the initialization distribution of the last layer of SIREN, i.e., $W^{(L)}\sim\mathcal{U}\left(-\sqrt{6}/\left(n^{3/4}_{L-1}\omega\right),\sqrt{6}/\left(n^{3/4}_{L-1}\omega\right)\right)$, where $n_{L-1}$ denotes the width of the (L1)(L-1)-th layer.
Sinc NF (Saratchandran et al., 2024c): In this case, we use activation function as a sinc function, i.e., $\sigma(Z)=\sin(\omega x)/\omega x$, where $\omega$ is the fixed parameter. We use $\omega=8$ for all experiments, except for audio fitting (in this case, we use $\omega=16$).
Xavier uniform initialization (Glorot and Bengio, 2010): In this case, we use the standard Xavier uniform initialization for the SNF architecture, i.e., $W^{(l)}\sim\mathcal{U}\left(-\sqrt{6}/\sqrt{n_{l-1}+n_{l}},\sqrt{6}/\sqrt{n_{l-1}+n_{l}}\right),\quad l\in[1,L]$. We note that the Xavier uniform initialization in SNF does not yield reasonable results. However, the purpose of presenting it is to highlight the initialization principle of SNFs (Appendix A), which guarantees stable training.
@@
F.5 Neural dataset experiments
@@
[…]
F.6 Neural radiance fields experiments
In this subsection, we provide details and quantitative results of novel view synthesis via neural radiance fields (NeRFs) (Mildenhall et al., 2020). Neural radiance fields (NeRFs) represent scenes using neural fields. More specifically, NeRF employs neural fields (i.e., MLPs) with finite amount of training data to capture the scene, aiming to recover the scene in a continuous manner. At the end, we input unseen coordinates into the model and evaluate the test PSNR.
Datasets. We use the ‘Lego’ and ‘Drums’ data from the ‘NeRF-synthetic’ dataset (Mildenhall et al., 2020), which is publicly available online. Each dataset contains 100 training images, 100 validation images, and 200 test images, along with their corresponding camera directions.
Implementation details. We mainly follow the settings of WIRE (Saragadam et al., 2023), using the ‘torch-ngp’ codebase. Specifically, we use a 4-layer volume density network and a 4-layer color network, each with a width of 182, except for WIRE. This exception is due to the complex-valued weights of WIRE, where we use a width of 128 for fair comparison. For the WS network, we apply weight scaling only to the volume density network.
Task 1: training with $\times800$ images. In this case, we use $\omega_{0}=15$ and $\omega_{h}=5$ for the SNF family, frequency lifting factor $k=10$ for ReLU + P.E., $s=20$ for GaussNet, and $\omega=5$, $s=10$ for WIRE. Additionally, we use positional encoding scheme only for ReLU networks. We use learning rate 1e021e-02 for ReLU+P.E., 1e031e-03 for FFN, 3e043e-04 for SNF family, 1e031e-03 for GaussNet, and 6e046e-04 for WIRE. We train all models until 300 epochs. We train the model with 100 images and evaluate the test PSNR with 100 untrained images.
Task 2: training with $\times200$ images, with longer epochs. Additionally, as in Saragadam et al. (2023), we train model with $200\times 200$ images, and trained until further epochs (600 epochs). In this case, we use $\omega_{0}=15$ and $\omega_{h}=3$ for the SNF family, frequency lifting factor $k=10$ for ReLU + P.E., $s=20$ for GaussNet, and $\omega=20$, $s=40$ for WIRE. Also, we use positional encoding scheme only for ReLU networks. We use learning rate 6e036e-03 for ReLU + P.E., 3e033e-03 for FFN, 9e049e-04 for SNF family, 1e031e-03 for GaussNet, and 9e049e-04 for WIRE. We train the model with 100 images and evaluate the test PSNR with 100 untrained images.
Results. We report the test PSNR and learning curve in Table 4 and Fig. 16 respectively. For both datasets and both training resolutions, WS achieves a higher test PSNR compared to other baselines. We emphasize that WS primarily focuses on fast training of neural fields, but here we also demonstrate its good generalization quality. Qualitative results can be found in Fig. 21.
Method $\times800$ (300 Epochs) $\times200$ (600 Epochs)
Lego Drums Lego Drums
ReLU + P.E. (Mildenhall et al., 2020) 26.02 20.97 31.57 26.21
SIREN (Sitzmann et al., 2020b) 27.68 24.05 32.02 27.73
FFN (Tancik et al., 2020) 26.96 22.10 30.40 24.68
GaussNet (Ramasinghe and Lucey, 2022) 25.95 22.42 30.61 26.24
WIRE (Saragadam et al., 2023) 26.71 23.62 32.09 27.73
NFSL (Saratchandran et al., 2024b) 27.79 24.04 32.10 27.64
Weight Scaling (ours) 28.17 24.10 32.48 27.86
Table 4: Quantitative results: test PSNR for NeRF experiments.
(a) Lego (800 $\times$ 800 images, 300 epochs)
(b) Drums (800 $\times$ 800 images, 300 epochs)
(c) Lego (200 $\times$ 200 images, 600 epochs)
(d) Drums (200 $\times$ 200 images, 600 epochs)
Figure 16: Test PSNR curves for NeRF experiments under different settings. The first row represents experiments conducted with 800 $\times$ 800 images over 300 epochs, while the second row shows experiments conducted with 200 $\times$ 200 images over 600 epochs. WS initialization consistently achieves better test PSNR compared to the baselines.
F.7 Solving differential equations with partial observations
In this subsection, we provide details about solving wave equation: a type of second order time-dependent partial differential equation (PDE) describing wave propagation in time $t$. The initial conditions are only the given information, neural fields aim to extrapolate the equation in timescale domain. We use the same formulation as in Sitzmann et al. (2020b).
Datasets and training details. We set $f(x)$ (i.e., neural network) to reconstruct and extrapolate Gaussian pulse wave equation, same setup as in Sitzmann et al. (2020b). Input coordinates are 3-dimensional (i.e., 2-dimensional spatial and 1-dimensional temporal timescale from $t=0$ to $4$), and outputs 1-dimensional Gaussian pulse value for each spatiotemporal coordinate. Other hyperparameters are set the same as in Sitzmann et al. (2020b), except for the number of iterations (we trained for up to 40k iterations).
Results. We present the qualitative results in Fig. 22. However, in Fig. 22, the WS network consistently predicts the boundary of the Gaussian, compared to other architectures [^4] Since the ground truth Gaussian pulses for $t>0$ are unavailable, they can only be reconstructed using a PDE solver or utilizing finite element method.. Qualitative comparisons are provided using coordinate-wise error (i.e., MSE) at $t=0$. It can be seen that WS-SNF not only faithfully reconstructs the given observation but is also able to generalize accurately to further timesteps. For WIRE and MFN, they failed to reconstruct the pulse; therefore, we do not provide the results.
F.8 Qualitative results
@@
Ground truth
ReLU + P.E.
FFN
SIREN
MFN
GaussNet
WIRE
NFSL
Sinc NF
WS (ours)
@@
ReLU + P.E.
FFN
SIREN
Ground Truth
MFN
GaussNet
WIRE
NFSL
Sinc NF
WS (ours)
@@
Ground truth
ReLU + P.E.
FFN
SIREN
MFN
GaussNet
WIRE
NFSL
Sinc NF
WS (ours)
@@
Ground truth
ReLU + P.E.
FFN
SIREN
MFN
GaussNet
WIRE
NFSL
Sinc NF
WS (ours)
@@
Lego (ground truth)
ReLU + P.E.
SIREN
GaussNet
FFN
WIRE
NFSL
WS (ours)
Drums (ground truth)
ReLU + P.E.
SIREN
GaussNet
FFN
WIRE
NFSL
WS (ours)
Figure 21: Novel view synthesis: qualitative results ($800\times 800$ images).
Ground truth ($t=0$, which is the only one given).
ReLU + P.E. (7.7e-03)
SIREN (1.8e-02)
GaussNet (4.3e-03)
NFSL (7.8e-04)
WS (2.8e-05)
Figure 22: Solving PDE from partial observation: qualitative results. Each column represents $t=$ 0, 0.07, 0.14, and 0.21 from left to right. The parentheses in the caption* represent the MSE at $t=0$.
F.9 Training for significantly larger iterations
In this section, we present the learning curves for Kodak image dataset over significantly longer iterations. We report the average train and test PSNR values for 5000 training steps, in which most methods converge to saturation. The learning rate scheduler is employed, resulting in improved performance across most methods under extended training setting, and the learning rates are tuned for each architecture to work well with a scheduler.
(a) Train PSNR curves for Kodak dataset
(b) Test PSNR curves for Kodak dataset
Figure 23: PSNR curves over 5000 iterations. We compared the PSNR of weight scaling to other baselines.
Weight scaling demonstrates superior train accuracy, achieving higher PSNR with substantial acceleration. Moreover, even at saturation, weight scaling outperforms other methods, with the highest PSNR. Notably, our methods maintain its generalization ability even for extended training. Weight scaling reaches its peak test PSNR (37.45 dB), which is comparable to or exceeds that of other methods, much faster, demonstrating the fastest convergence in both train and test performance.
We utilize learning rate 1e021e-02 for Xavier initialized SNF, 1e031e-03 for ReLU + P.E., 1e031e-03 for FFN, 1e041e-04 for SNF family, 1e031e-03 for GaussNet, 1e021e-02 for MFN, and 1e041e-04 for WIRE, with Adam optimizer.